### v1

@@ 4.1 Temporal analysis of annotations
Figure 3: Distribution of annotation times with and without pre-annotations. Bin steps of 20 seconds were employed. The kernel density estimation plot illustrates the trend for both distributions.
Table 1: Analysis of the independence between two sets through parametric statistical testing. The outcomes of the statistical test for independence are detailed in the upper section. The lower section provides a statistical overview of the two distributions.
Parametric test Statistic $p$-value Satisfies independence assumption ($p$ << 0.005)?
$t$-test -3.6391 0.0004 True
One-way ANOVA 13.2428 0.0004 True
Wilcoxon signed-rank 853.0000 0.0002 True
Annotation time
Type of annotation #Annotations #Videos Mean (s) STD (s)
With pre-annotation 90 30 146.42 79.28
Without pre-annotation 90 30 190.44 72.63
Comparing annotation times per video is less meaningful than per annotator since human factors exert the most significant influence. Figure 4 illustrates that 72% of annotators took less time to complete their annotations with pre-annotation, with a median improvement of 35%. While acknowledging potential influences from human factors and video content, efforts have been made to isolate annotation times as much as possible. Figure 5 explores the relationship between annotation times, the number of classes per video, and their duration, revealing no direct correlation.
@@
Figure 4: Comparison of annotation times per anonymized user. Each point corresponds to the annotation times of an annotator, and the median of all annotation times with and without pre-annotation has been taken. Figure 4(a) contains the total annotation times per video, while in figure 4(b), these times are normalized by the duration of each video. Points above the green area indicate that the median annotation time with pre-annotation is shorter than without, whereas points above the red area suggest the opposite.
Figure 5: Figure 5(a) displays the median annotation time for all videos grouped by the number of non-unique classes. Figure 5(b) is the median annotation time per video duration, with intervals set at 5 seconds. In both cases, the data is separated by annotation type.
Lastly, a phenomenon extracted from the data is the impact of training. According to annotator feedback, pre-annotation utilization improved with experience in annotation. We removed the first four annotations from each annotator to observe these effects and compared the results with those obtained in Figure 4. Figure 6 highlights the vectors depicting the differences. Most annotators benefited more from pre-annotations in their last six out of ten.
Figure 6: Comparison of results obtained in Figure 4, excluding the first four annotations. Each red arrow represents the change for each annotator. The farther from the boundary between the zones, the greater the increase in annotation time difference with and without pre-annotation.
4.2 Quality Analysis of Annotations
The quality analysis has two components: we attempt to estimate the ground truth through our aggregation methodology, and the second involves comparing the annotations among themselves. The intrinsic freedom of the test does not allow for a broad consensus on what is considered a “correct annotation”.
When aggregating such annotations, it is crucial to consider that labels are independent of absolute values, and a permutation of labels should not alter the result. Only the clustering of frames is valued. This consideration is essential both during aggregation and when selecting evaluation metrics. The methodology for obtaining our ground truth is as follows:
  • Extract all annotations of the same type from each frame.
  • Create a tuple with the annotations and hash it. Using hashing in this context helps create a unique identifier for each different group of frames without relying on the specific values of the labels. This approach allows us to handle cases with an infinite number of classes.
  • The new labels are the hashes of each group, considered as the ground truth. This way, we preserve only the parts where annotations coincide, even in the presence of noisy annotations.
  • Calculate metrics tailored to our use case with the ground truth and each annotation, averaging all combinations.
This approach helps address the noise issue and enables the study of patterns of similarity in clustering. Considering the hashes of frame annotations as the ground truth results in clusters where all annotations have deemed these frames to be grouped similarly. Therefore, we can consider these clusters as “strong clusters”, while the rest of the information is discarded.
The next step involves selecting metrics that measure only the shared information between the two classifications, disregarding the absolute value of the label, permutation of classes, and the number of classes. The selected metric is Adjusted Mutual Info (Adjusted, Normalized and V-measure) defined by the equation:
$\text{Mutual Information}=\sum_{i=1}^{\mathbf{|U|}}\sum_{j=1}^{\mathbf{|V|}}\frac{\mathbf{|U_{i}\cap V_{j}|}}{\mathbf{N}}\log\left(\frac{\mathbf{N|U_{i}\cap V_{j}|}}{\mathbf{|U_{i}||V_{j}|}}\right).$ (1)
Here, $\mathbf{|U_{i}|}$ represents the number of samples in cluster $\mathbf{U_{i}}$, $\mathbf{|V_{j}|}$ represents the number of samples in cluster $\mathbf{V_{j}}$, and $\mathbf{N}$ is the total number of samples. Table 2 presents the results, indicating no significant difference in the classification quality.
Table 2: Statistics from the test comparing annotations of different types with the aggregated ground truth.
Type of Annotation Adjusted Mutual Information Normalized Mutual Info V-measure
With pre-annotation 0.64 0.64 0.64
Without pre-annotation 0.63 0.64 0.64
Another aspect to explore is whether a discernible trend toward annotation homogenization exists when utilizing pre-annotation. Since a lot of freedom was given in annotations precisely to deduce the effect of pre-annotation, Table 3 contains the results of our homogenization test. Metrics were calculated with all combinations of pairs of annotations of the same type for each video, and the mean of all these combinations is represented. As observed in the table, the similarity between annotations with pre-annotation is higher, indicating greater homogeneity. This capability empowers designers to steer outcomes by leveraging pre-annotations to attain specific annotation nuances or uphold quality standards, thereby mitigating annotator fatigue.
Table 3: Statistics from the test on the homogeneity of annotations of different types.
Type of Annotation Adjusted Mutual Information Normalized Mutual Info V-measure
With pre-annotation 0.67 0.67 0.67
Without pre-annotation 0.62 0.62 0.62
4.3 Visual Similarity: Evaluating Results with CLIP
The CLIP [18] similarity (denoted as CLIP-SIM) metric has been widely employed in various studies [33, 23, 29, 17] to assess image generation models primarily. However, we aim to adapt this metric for video evaluation, following a methodology similar to FVD [25]. After encoding frames with the “Vit-32” image encoder from CLIP, we calculated the Silhouette Score using the cosine distance. This metric, in conjunction with embeddings and a distance function between them, aims to estimate the quality of clustering based on the semantic content of frames.
As illustrated in Figure 7, we observed a higher Silhouette Score in 25 out of 30 pre-annotated videos. The mean Silhouette Score with pre-annotation was 0.41, whereas without pre-annotation, it was 0.28, indicating a significant difference of 52.10%.
These results are expected since pre-annotation involves starting with a frame classification that maximizes the Silhouette Score. Considering this metric as an objective measure of classification quality, which respects the semantic content of the video frames, it can be inferred that pre-annotation enhances the annotation’s overall quality.
Figure 7: Comparison of the median Silhouette Score for all annotations of the same type within each video. Each point represents a video, totaling 30 points. Points in the green area indicate that the median Silhouette Score of pre-annotated annotations is higher than without annotation, and vice versa for the red area.
@@ 5 Conclusions
We have successfully implemented the Human-in-the-Loop paradigm for video annotation using automatic pre-annotations and conducted a full test to evaluate the annotation process, comparing its results with and without using these AI-boosted pre-annotations.
We observed that 70% of our novel annotators achieved a 35% reduction in their annotation time using the proposed HITL annotation scheme while delivering similar annotations.
We observed that the annotations differed, yet they remained equally valid. This divergence may relate to the freedom provided in the test, and even when analyzing the semantic content using the CLIP vision model, an improvement in classification concerning semantic content was noted. This suggests that introducing a bias that can be advantageous or detrimental depending on the test design. Moreover, it can be considered an additional factor in shaping the output of annotations, such as determining the level of detail or the type of actions to be classified.
Another valuable and qualitatively unmeasurable aspect throughout the process was the feedback from the annotators. They expressed that the pre-annotations reduced annotation fatigue, facilitating easier annotation with the tool. Such insights gain significance when extrapolated to routine annotation projects that entail hundreds of hours and involve a larger number of annotators.
Finally, the methods developed to conduct these experiments and the technology used in the pre-annotation have been devised to ease the annotation process. Nonetheless, the AI boost in pre-annotation may impact the workforce currently performing annotation duties, affecting the whole sector. First, the proposed technology may reduce dependence on human labor, ultimately relegating annotators to the most complex and challenging cases. Second, these technologies raise multiple concerns regarding the digital rights of the data and the annotations and how to preserve ownership and control it. From a technological standpoint, we understand that digital transformation will also affect the data annotation market. Evaluating the capabilities covered in this contribution sets the path for a deeper and broader discussion. More importantly, it will provide evidence of the pivotal role of the AI-powered HITL in transforming the data industries.
Acknowledgments
We appreciate the volunteer annotators for their dedication and time, as well as the Label Studio developers for providing an open-source annotation platform for this project.
Research framed under HADA project “Research on advanced data annotation tools" (2021/C005/00146323), funded by EU Next Generation (through Ministry of Economy and Digital Transformation & RED.ES).
Disclosure of Interests. Juan Gutiérrez received research funding at UPM from a project funded by Sigma Cognition.
Appendix 0.A Experimental evaluation CrowdWorkSheets
[As supplemental material, we include further details of the video annotation test that was conducted. This material should be published along with the manuscript and the results obtained.]
@@ Appendix 0.A Experimental evaluation CrowdWorkSheets
To further describe our annotation process, we include our responses to the CrowdWorkSheets questionnaire [6]. CrowdWorkSheets is a framework for transparent and ethical dataset annotation that emphasizes the annotation process’s documentation, considering the annotators’ identities and experiences. The framework provides guidance on ethical considerations, annotator selection, and documentation of the annotation process, aiming to address unresolved ethical issues in crowdsourcing. More importantly, it allows us to share a complete summary of the annotation process conducted to ease replicating the test and to highlight the key aspects behind its execution.
0.A.1 Task Formulation
@@
  • We assume that annotators are committed to conducting the test accurately. This assumption is based on their status as trusted individuals within a university department and their academic interest in the test, given their involvement in projects of the same theme.
@@
Annotators will be tasked with annotating crime videos captured by surveillance cameras. The estimated duration of the test is less than 30 minutes, with 2-4 minutes allocated for each video. Prior filtering has been conducted to exclude videos containing overly sensitive imagery.

### v2

@@ 4.3 Temporal Analysis: Efficiency Gains
Figure 5: Distribution of annotation times with and without Pre-Annotations. We used bin steps of 20 seconds. The kernel density estimation plot illustrates the trend for both distributions.
Table 2: Statistical comparison of annotation times between assisted and unassisted workflows. Both parametric ($t$-test) and non-parametric (Wilcoxon) tests confirm a statistically significant difference between the distributions ($p<0.005$).
Test Statistic $p$-value
$t$-test -3.64 0.0004
Wilcoxon 853.0 0.0002
Table 3: Descriptive statistics for annotation time (in seconds).
Condition Mean STD
Assisted 146.42 79.28
Unassisted 190.44 72.63
As shown in Table 3, the results indicate a mean time reduction of 23.11% across the cohort. However, when normalizing for individual annotator speed, we observe that 72% of participants worked faster with the automatic suggestions, achieving a median improvement of 35%. Figures 7a–b break this down per anonymized user; points above the green area indicate efficiency gains. Notably, we found no correlation between the time savings and the video duration or class count (Figure 6), suggesting the efficiency gain is robust to varying content complexity.
A temporal analysis of the user sessions (Figures 7c–d) reveals that the benefit of Pre-Annotation increased over time. By excluding the first four “training” videos, the efficiency gap widens, indicating that as users become familiar with the tool, they become more adept at verifying suggestions rather than re-doing them.
@@
(a) Impact of Class Frequency
(b) Impact of Video Duration
Figure 6: Efficiency gains against content complexity. The plots display median annotation times stratified by (a) non-unique classes per video and (b) video duration (5-second bins).
(a) Total Time
(b) Normalized Time
(c) Raw Shift
(d) Normalized Shift
Figure 7: Per-user efficiency analysis. (a–b) Total and normalized annotation time per user; points in the green zone indicate users who were faster with Pre-Annotation assistance (72% of participants). (c–d) Changes in efficiency after excluding the first four training videos; each red arrow represents the change for an annotator.
4.4 Quality Analysis: Convergence and Inter-Annotator Stability
Defining the exact temporal boundaries of anomalous events involves inherent subjectivity. Rather than enforcing rigid, artificial constraints, we evaluate how the workflow manages this natural ambiguity through two dimensions: alignment with a consensus baseline (Accuracy) and agreement among annotators (Homogeneity).
We establish a Consensus Baseline by aggregating the classifications of all six annotators per video. By grouping identical frame-level label permutations, we isolate the “strong clusters” where the majority of users agree, filtering out individual noise. We then quantify performance using Adjusted Mutual Information (AMI). An AMI of 1.0 indicates perfect alignment, while 0.0 implies random correlation.
Table 4 presents the comparative analysis. To address concerns that assistance might induce blind acceptance, a passive anchoring bias in which users accept incorrect suggestions, we examine the alignment of individual sessions against the Consensus Baseline. If the Pre-Annotations introduce systematic biases or if users were passively accepting hallucinated labels, we would expect a degradation in alignment with this baseline. However, both the assisted and unassisted groups maintain a nearly identical alignment score (AMI $\approx$ 0.64). This parity indicates that Pre-Annotation suggestions do not distort user judgment or induce deviation from collective semantic consensus. Instead, users actively verify and correct automatic suggestions, preserving alignment with the natural group wisdom observed in the unassisted condition.
We evaluated Inter-Annotator Consistency (Homogeneity). Here, a significant divergence appears (Table 4, down rows). The assisted group demonstrated notably higher agreement with each other (AMI: 0.67) compared to the unassisted group (AMI: 0.62).
In the absence of strict definitions, unassisted annotators exhibited high variance in boundary selection (“jitter”). The Pre-Annotation workflow mitigated this not by enforcing an arbitrary correct answer (which would lower the consensus score), but by providing a “semantic anchor.” This proves that the workflow successfully acts as a standardization mechanism, reducing the stochastic noise inherent to human labeling without altering the fundamental semantic classification of the events.
Table 4: Comparative analysis of annotation metrics. We evaluate two dimensions: alignment with the Consensus Baseline (Accuracy) and Inter-Annotator Consistency (Homogeneity). Abbreviations: Adjusted Mutual Information (AMI), Normalized Mutual Information (NMI), and V-measure (V).
Condition AMI NMI V
Consensus Alignment (Accuracy)
Assisted 0.64 0.64 0.64
Unassisted 0.63 0.64 0.64
Inter-Annotator Consistency
Assisted 0.67 0.67 0.67
Unassisted 0.62 0.62 0.62
4.5 Semantic Standardization: Latent Space Analysis
We evaluate the structural coherence of the annotations using latent-space analysis. We compute the Silhouette Score of the annotated segments using the CLIP image encoder (Radford et al., 2021) and cosine distance. While evaluating with the same encoder family used for generation introduces inductive bias, we employ this metric specifically to quantify standardization, not for ground truth validation. In high-ambiguity domains, unassisted human annotation is prone to significant stochastic variance (semantic jitter); we hypothesize that the assisted workflow acts as a regularizer.
We adopt the Silhouette Score—typically used for generative fidelity (Zheng et al., 2023; Tao et al., 2023)—as a proxy for internal consistency (Shahapure and Nicholas, 2020). As illustrated in Figure 8, the assisted workflow improved structural coherence in 25 of 30 videos. The mean Silhouette Score increased from 0.28 (unassisted) to 0.41 (assisted), a relative improvement of 52.10%.
Crucially, this increased standardization does not degrade semantic accuracy. As reported in Table 4, the assisted cohort maintained a Consensus Alignment (AMI) of 0.64, equivalent to the unassisted baseline. Consequently, the increase in Silhouette Score indicates that the Pre-Annotation pipeline successfully filters out inter-annotator noise without distorting the underlying semantic consensus.
Figure 8: Latent Space Standardization. Each point represents a video. Points in the green region denote videos where the assisted workflow yielded higher Silhouette Scores, indicating a reduction in stochastic annotator variance (jitter) relative to the unassisted baseline.
@@ 5 Conclusions
In this work, we introduced a Human-in-the-Loop (HITL) workflow for dense temporal video segmentation that reduces the cost of producing high-quality labels by reframing annotation as verification-and-editing rather than from-scratch creation. We move beyond simulated studies and evaluate the workflow in a controlled experiment with 18 annotators labeling open-world footage, allowing us to measure how AI assistance changes the annotation process in practice.
Our results show that model-generated Pre-Annotations substantially improve efficiency: annotation time decreases by a median of 35% for most participants. Importantly, these gains are not achieved simply by making annotators agree more at the expense of meaning. Instead, the resulting segmentations exhibit greater structural standardization and greater cross-annotator homogeneity, consistent with algorithmic guidance that helps annotators converge on clearer temporal boundaries rather than introducing semantic drift.
To support reproducibility and enable follow-up work on human–model interaction dynamics, we release the full open-source implementation of the pipeline, the analysis code, and an anonymized dataset of interaction traces. Together, these artifacts provide a validated methodology for studying and benchmarking AI-assisted annotation and offer a practical foundation for scaling the creation of complex, densely labeled video datasets.
Impact Statement
This work advances data-centric AI by improving Human-in-the-Loop workflows for dense temporal video segmentation, reducing the cost and effort required to build high-quality datasets for video understanding and anomaly detection. By shifting the role of the annotator from creating labels from scratch to reviewing and correcting model-generated Pre-Annotations, the workflow can increase efficiency; but it also changes the human–AI interaction in ways that matter for data quality. In particular, Pre-Annotations can introduce anchoring bias, leading annotators to unintentionally be steered toward the model’s proposed boundaries or categories. Effective systems must therefore be designed to preserve human agency and encourage active verification rather than passive acceptance of algorithmic outputs.
Finally, these workflows are often used on open-world footage that may include sensitive, explicit, or distressing content. Researchers and practitioners should apply rigorous ethics and safety practices—such as content warnings, opt-out mechanisms, and appropriate support—to protect the’ psychological well-being of annotators.
Appendix A Video Encoder Evaluation
In this appendix, we provide a detailed evaluation of the cross-modal video encoder used in our Pre-Annotation Pipeline. While the main body of this paper focuses on the human-centric evaluation of the workflow, verifying the robustness of the underlying model is essential to validate the quality of the suggestions provided to the annotators.
A.1 Evaluation Protocols
Video-Text Matching. We evaluated the cross-modal encoder’s ability to pair videos with corresponding text snippets accurately. The model predicts the most probable (video, text) pairs based on embeddings generated by the encoders. These embeddings, representing the semantic essence of both modalities, are compared using cosine similarity, scaled by a temperature parameter, and normalized via softmax to produce a probabilistic prediction of the correct pairing.
The model was trained for 30 epochs, balancing time and learning performance, with monitoring to select the best-performing epoch for final evaluation. All videos were divided and matched with the corresponding descriptions provided in the UCA dataset. During the accuracy calculation, we introduced randomness by selecting video-clips from within the same video and across different videos.
[…]
A.2 Quantitative Results and Ablation
Training Results. Table 5 reveals that the ViT-B/16 architecture with a batch size of 10 and 16 sampled frames achieved the highest top-1 accuracy of 89.68% and a top-5 accuracy of 99.64%. Other configurations, such as the ViT-B/32 with varying batch sizes and sampled frames, also demonstrated competitive performance, though with some variation in accuracy depending on the specific setup.
Effect of Frame Sampling. Frame sampling is critical in determining the balance between computational efficiency and retrieval accuracy. We reduce the computational load by selecting a subset of frames rather than processing the entire video. However, with scarce sampling frames, one may omit important temporal or semantic cues, degrading the model’s performance.
We experimented with different sampling intervals to evaluate the effect of frame sampling, as shown in Table 5. When sampling fewer frames (for example, 5 or 16 frames), the retrieval accuracy improves significantly, with Acc@1 reaching 89.68% for ViT-B/16 with 16 sampled frames. Sampling 30 frames with ViT-B/16 slightly decreases the Acc@1 score to 73.59%, indicating a diminishing return when too many frames are processed simultaneously. This suggests that there is an optimal range of frame sampling that balances both efficiency and performance. Sampling too few frames may hinder accuracy while sampling too many frames may introduce unnecessary redundancy and increase computational complexity.
Impact of Model Backbone. The model architecture, particularly the backbone choice, is crucial in retrieval performance. We compared two versions of the CLIP architecture: ViT-B/16 and ViT-B/32. Table 5 details that the ViT-B/16 backbone consistently outperforms ViT-B/32 in accuracy. For example, with a batch size of 10 and 16 sampled frames, ViT-B/16 achieves 89.68% Acc@1 and 99.64% Acc@5, significantly outperforming ViT-B/32, which achieves Acc@1 scores of around 73%.
The improved performance of ViT-B/16 is likely due to its ability to capture more detailed spatiotemporal relationships, thanks to its higher-resolution feature representations. While ViT-B/32 provides competitive results and is less computationally expensive, the added granularity of ViT-B/16 proves more beneficial for tasks requiring precise retrieval across long and complex video sequences.
Table 5: Performance comparison of various CLIP fine-tuning configurations. The table shows results for different backbone architectures (ViT-B/16 and ViT-B/32), batch sizes, and numbers of sampled frames. Performance metrics are evaluated using Acc@1 and Acc@5.
Backbone Batch Size Sampled Frames Acc@1 (%) Acc@5 (%)
Random - - $8.17$ $39.89$
ViT-B/16 10 16 89.68 99.64
30 30 73.59 95.30
ViT-B/32 5 30 73.71 95.20
30 5 73.71 95.20
20 32 81.76 97.75
100 40 59.59 86.24
A.3 Attention Shift
To understand how the fine-tuned encoder processes temporal dynamics, we analyzed the self-attention maps of the Vision Transformer (ViT) (Vaswani et al., 2017; Radford et al., 2021). Rather than treating the model as a black box, we visualized the attention weights to verify that the model focuses on relevant spatiotemporal actions.
Methodology. We extracted the self-attention weights from the multi-head attention layers of the ViT. Specifically, we computed the cumulative joint attention across layers and isolated the attention maps associated with the class (CLS) token. Since the CLS token serves as the global aggregate representation of the sequence, its attention map reveals which spatial patches the model deems most significant for the final classification. We averaged these weights across heads and upscaled them to match the input frame resolution for visualization (Li et al., 2023; Ma et al., 2022a).
Visual Analysis. Visualizing attention maps across various layers of the fine-tuned Vision Transformer (ViT) reveals essential insights into the model’s interpretation and processing of video data. Figure 9 illustrates the model’s ability to focus on relevant spatial and temporal aspects within input sequences, reflecting an enhanced understanding of inter-object relationships and scene dynamics.
The analysis of these attention maps reveals several key insights. In the “Vandalism034” video-record, the model effectively tracks the aggressive interactions between individuals and the trash can, highlighting the critical movements that define the event. Similarly, in the “Stealing004” case, the model accurately identifies the coordinated actions of the individuals involved in the theft, with a strong focus on the movement of the motorcycle and the interactions surrounding it. These examples showcase the model’s capability to discern complex, temporally extended activities.
In the “Normal_Videos_352” video-record, which depicts routine activities like refueling cars, the attention maps demonstrate the model’s ability to recognize and focus on key actions, such as the departure of the car and the movements of individuals at the gas station. This indicates that the model can distinguish routine behaviors from more significant or unusual ones. In the “Abuse005” video-record, involving a prison guard’s interactions, the model focuses on the guard’s movements and posture, reflecting an understanding of the scene’s context and the key actions that convey the narrative.
These findings underscore the model’s inherent ability of the proposed encoder to learn and generalize temporal relationships through fine-tuning on video datasets, even without explicit temporal modeling modifications. The attention mechanisms allow the ViT to prioritize important features and actions, leading to enhanced performance in video understanding tasks.
Refer to captionOriginalViT-B/32Trained
Vandalism034: “A group of people vandalizes a black trash can by repeatedly kicking, pushing, and hitting it, with one man particularly aggressive, using both his feet and elbow while holding a bottle.” Stealing004: “A man in a blue shirt directed a man in a striped shirt to move forward, then helped him push a black motorcycle out of a yard and onto the road.”
Refer to captionOriginalViT-B/32Trained
Normal_Videos_352: “A white car drives away, followed by a man in a purple shirt and a woman in a blue top refueling their cars and leaving to pay.” Abuse005: “A prison guard wearing black trousers, gray short-sleeved uniform and hat sat in front of a white table and looked at him, stood up.”
Figure 9: Visualization of attention maps on UCF-Crime. Each merged image contains three stacked rows (Original, ViT-B/32, Trained) and two video samples per row (left / right). Rotated labels are placed at the centers of the three thirds of the merged image height; descriptions are side-by-side beneath each merged image.
Appendix B Experimental evaluation CrowdWorkSheets
@@
To further describe our annotation process, we include our responses to the CrowdWorkSheets questionnaire (Díaz et al., 2022). CrowdWorkSheets is a framework for transparent and ethical dataset annotation that emphasizes the annotation process’s documentation, considering the annotators’ identities and experiences. The framework provides guidance on ethical considerations, annotator selection, and documentation of the annotation process, aiming to address unresolved ethical issues in crowdsourcing. More importantly, it allows us to share a complete summary of the annotation process conducted to ease replicating the test and to highlight the key aspects behind its execution.
B.1 Task Formulation
@@
  • We assume that the annotators are committed to conducting the test accurately. This assumption is based on their status as trusted individuals within a university department and their academic interest in the test, given their involvement in projects of the same theme.
@@
Annotators will be tasked with annotating open-world videos. The estimated duration of the test is less than 30 minutes, with 2-4 minutes allocated for each video. Prior filtering has been conducted to exclude videos containing overly sensitive imagery.